# Coupler Position Optimization and Channel Estimation for Flexible Coupler Aided Multiuser Communication

### v1

@@ V-B Channel estimation performance
Next, simulations are conducted to validate the effectiveness of the proposed distributed and centralized FC channel estimation algorithms. The normalized mean square error (NMSE) results are obtained by averaging over all users and $500$ Monte Carlo trials.
@@
In Fig. 8, we plot the NMSE versus the average received SNR for the proposed centralized FC channel estimation with different numbers of pilot blocks $V$. We adopt exhaustive measurement as a high-overhead benchmark scheme, where the coupler movement region is discretized into a large number ($D=400$) of candidate locations and the effective channel is sequentially measured over all candidates, resulting in on the order of $MND$ measurements per pilot block. From Fig. 8, it can be seen that increasing $V$ significantly improves the NMSE performance of all schemes, especially in the low to medium SNR regime, since varying the coupler positions across blocks generates more diversified effective channel observations and enhances angular sparse reconstruction through stacked measurements. Moreover, as observed from Fig. 8, exhaustive measurement becomes more competitive at high SNR due to its dense direct sampling with high computational complexity. In contrast, the proposed FC method with a larger $V$ achieves lower NMSE at low SNR, since joint sparse reconstruction across blocks provides an inherent noise-averaging effect while requiring substantially fewer measurements.
@@
In Fig. 9, we plot the NMSE versus the pilot length $\tau$ for the proposed distributed FC channel estimation and exhaustive measurement, under two coupler numbers $N\in\{2,6\}$. It is observed that increasing $\tau$ improves the estimation accuracy for all schemes, since longer pilot sequences enhance the reliability of the pilot correlation outputs. In the short pilot regime, the proposed distributed method achieves lower NMSE than exhaustive measurement for both values of $N$, as it exploits structured sparsity and performs fusion using low-dimensional sufficient statistics, which provides effective noise suppression when pilots are limited. As $\tau$ increases, exhaustive measurement becomes more competitive due to its dense direct sampling. Moreover, increasing $N$ leads to higher NMSE for both schemes, indicating a higher pilot requirement to maintain the same estimation accuracy.
@@
In Fig. 10, we compare the NMSE performance of the proposed centralized and distributed FC channel estimation schemes versus the average receive SNR. It is observed that both schemes achieve lower NMSE as the SNR increases, since pilot correlation becomes more reliable and stacked observations across pilot blocks provide more informative measurements for sparse reconstruction. Moreover, the centralized scheme consistently outperforms the distributed one over the entire SNR range, as it can jointly exploit the sparsity structure across all antennas and users, whereas the distributed scheme relies on local processing at FC antennas with limited information fusion. The performance gap becomes more pronounced at medium and high SNR, where the estimation error is dominated by information loss due to decentralization rather than noise. Nevertheless, the distributed FC channel estimation offers a favorable accuracy-complexity tradeoff, since it performs local processing and exchanges only low-dimensional sufficient statistics rather than raw pilot observations, thereby significantly reducing the communication signaling and computational burden at the CPU.
@@ VI Conclusions
We have proposed a distributed FC antenna array that enables mechanical beamforming through passive coupler repositioning around fixed active antennas, supported by a distributed architecture with local LPUs and a CPU. We have also investigated pilot-assisted centralized and distributed channel estimation solutions for the FC array. Simulation results validate the performance of the proposed FC array and channel estimation algorithms compared with benchmark schemes. In the proposed FC array, reconfigurable passive coupler positioning provides a cost-effective way to enhance communication performance without the need to add more RF chains. Moreover, the compact form factor and small footprint of the FC array make it particularly appealing for devices with stringent size, weight, and power (SWAP) constraints. For the future work, we will apply the proposed FC array to support diverse applications and services, such as being mounted on unmanned aerial vehicles (UAVs) with limited payload in the low-altitude economy. Beyond communications, we will investigate FC-enabled target sensing to achieve a more favorable sensing-communication tradeoff in integrated sensing and communication (ISAC) systems.

### v2

@@ V-B Channel estimation performance
Next, simulations are conducted to validate the effectiveness of the proposed distributed and centralized FCA channel estimation algorithms. The normalized mean square error (NMSE) results are obtained by averaging over all users and $500$ Monte Carlo trials.
@@
In Fig. 8, we plot the NMSE versus the average received SNR for the proposed centralized FCA channel estimation with different numbers of pilot blocks $V$. We adopt exhaustive measurement as a high-overhead benchmark scheme, where the coupler movement region is discretized into a large number ($D=400$) of candidate locations and the effective channel is sequentially measured over all candidates, resulting in on the order of $MND$ measurements per pilot block. From Fig. 8, it can be seen that increasing $V$ significantly improves the NMSE performance of all schemes, especially in the low to medium SNR regime, since varying the coupler positions across blocks generates more diversified effective channel observations and enhances angular sparse reconstruction through stacked measurements. Moreover, as observed from Fig. 8, exhaustive measurement becomes more competitive at high SNR due to its dense direct sampling with high computational complexity. In contrast, the proposed FCA method with a larger $V$ achieves lower NMSE at low SNR, since joint sparse reconstruction across blocks provides an inherent noise-averaging effect while requiring substantially fewer measurements.
@@
In Fig. 9, we plot the NMSE versus the pilot length $\tau$ for the proposed distributed FCA channel estimation and exhaustive measurement, under two coupler numbers $N\in\{2,6\}$. It is observed that increasing $\tau$ improves the estimation accuracy for all schemes, since longer pilot sequences enhance the reliability of the pilot correlation outputs. In the short pilot regime, the proposed distributed method achieves lower NMSE than exhaustive measurement for both values of $N$, as it exploits structured sparsity and performs fusion using low-dimensional sufficient statistics, which provides effective noise suppression when pilots are limited. As $\tau$ increases, exhaustive measurement becomes more competitive due to its dense direct sampling. Moreover, increasing $N$ leads to higher NMSE for both schemes, indicating a higher pilot requirement to maintain the same estimation accuracy.
@@
In Fig. 10, we compare the NMSE performance of the proposed centralized and distributed FCA channel estimation schemes versus the average receive SNR. It is observed that both schemes achieve lower NMSE as the SNR increases, since pilot correlation becomes more reliable and stacked observations across pilot blocks provide more informative measurements for sparse reconstruction. Moreover, the centralized scheme consistently outperforms the distributed one over the entire SNR range, as it can jointly exploit the sparsity structure across all antennas and users, whereas the distributed scheme relies on local processing at FCAs with limited information fusion. The performance gap becomes more pronounced at medium and high SNR, where the estimation error is dominated by information loss due to decentralization rather than noise. Nevertheless, the distributed FCA channel estimation offers a favorable accuracy-complexity tradeoff, since it performs local processing and exchanges only low-dimensional sufficient statistics rather than raw pilot observations, thereby significantly reducing the communication signaling and computational burden at the CPU.
@@ VI Conclusions
We have proposed a distributed FCA array that enables mechanical beamforming through passive coupler repositioning around fixed active antennas, supported by a distributed architecture with local LPUs and a CPU. We have also investigated pilot-assisted centralized and distributed channel estimation solutions for the FCA array. Simulation results validate the performance of the proposed FCA array and channel estimation algorithms compared with benchmark schemes. In the proposed FCA array, reconfigurable passive coupler positioning provides a cost-effective way to enhance communication performance without the need to add more RF chains. Moreover, the compact form factor and small footprint of the FCA array make it particularly appealing for devices with stringent size, weight, and power (SWAP) constraints. For the future work, we will apply the proposed FCA array to support diverse applications and services, such as being mounted on unmanned aerial vehicles (UAVs) with limited payload in the low-altitude economy. Beyond communications, we will investigate FCA-enabled target sensing to achieve a more favorable sensing-communication tradeoff in integrated sensing and communication (ISAC) systems.